The perimeter of a swimming pool is 96m. The width is 3m more than half the length of the pool. Find the length and breadth of the pool.
Answer by checkley77(12844)

P=2L+2W
W=L/2+3
96=2L+2(L/2+3)
96=2L+L+6
3L=96-6
3L=90
L=90/3=30 ans. for the length.
W=30/2+315+3=18 ans. for the length.
Proof:
96=2*30+2*18
96=60+36
96=96